8(b) For 25 army personnels, line of regression of weight of kidneys (Y) on weight of 

heart (X ) is Y = .0 399X + .6 934 and the line of regression of weight of heart on 

weight of kidney is X − .1 212Y + .2 461= .0 Find the correlation coefficient between 

X and Y and their mean values.


1
Expert's answer
2021-03-29T10:57:56-0400

Rewrite the regression equations uniformly


"Y = .0 399X + .6 934"

"X = .1 212Y - .2 461"


Denote mean values as "\overline{x}" and "\overline{y}", correlation coefficient as "r", standard deviations as "s_x" and "s_y". Then the basic formulas are


"Y=r\dfrac{s_y}{s_x}X-r\dfrac{s_y}{s_x}\overline{x}+\overline{y}"


"X=r\dfrac{s_x}{s_y}Y-r\dfrac{s_x}{s_y}\overline{y}+\overline{x}"


So find correlation coefficient from following


"\left\{\begin{array}{rr}r\dfrac{s_y}{s_x}=.0 399\\ \\ r\dfrac{s_x}{s_y}= .1 212\end{array}\right."


Multiply the equations


"r\dfrac{s_y}{s_x}\cdot r\dfrac{s_x}{s_y}=.0 399 \cdot .1 212" "\Rightarrow" "r^2=.00483588"


Thus

"r=\sqrt{.00483588}=.069540492"


Further, find mean values from


"-r\dfrac{s_y}{s_x}\overline{x}+\overline{y}= .6 934"


"-r\dfrac{s_x}{s_y}\overline{y}+\overline{x}=- .2 461"


or


"- .0 399\overline{x}+\overline{y}= .6 934"

"-.1 212\overline{y}+\overline{x}=- .2 461"


Construct the system of linear equations and solve it using Cramer's rule


"\left\{\begin{array}{rrr}- .0 399\ \overline{x}&+\ \overline{y}=& .6 934\\
\overline{x}&-.1 212\ \overline{y}=&- .2 461\end{array}\right."


"\det(A)= \begin{vmatrix}
 - .0 399 & 1 \\
 1 & -.1 212
\end{vmatrix}=" "-.0 399\cdot( -.1 212)-1\cdot1=" "-.99516412"


"\det(A_x)= \begin{vmatrix}
 .6 934 & 1 \\
 - .2 461 & -.1 212
\end{vmatrix}=" ".6934\cdot( -.1 212)-1\cdot(-.2461)=" ".16205992"


"\det(A_y)= \begin{vmatrix}
 - .0 399 & .6 934 \\
 1 & - .2 461
\end{vmatrix}=" "-.0 399\cdot( -.2461)-.6934\cdot1=" "-.68358061"


"\overline{x}=\dfrac{\det(A_x)}{\det(A)}=\dfrac{.16205992}{-.99516412}=-.162847431"


"\overline{y}=\dfrac{\det(A_y)}{\det(A)}=\dfrac{-.68358061}{-.99516412}=.686902388"


Answer


correlation coefficient

"r=.0695"


mean values

"\overline{x}=-.1628"

"\overline{y}=.6869"


Remark. Weight must be positive. Perhaps condition is incorrect.